OpenStudy (anonymous):

How Do u Do 5y+3=2y-3x+5

OpenStudy (anonymous):

3y=-3x+2........................y=-x+2/3

OpenStudy (anonymous):

Umm its a linear equationn So how Would i Balance it !/?/

OpenStudy (anonymous):

what do you mean by balance it?

OpenStudy (anonymous):

Yes i do

OpenStudy (precal):

Do you mean solve for y? \[5y+3=2y-3x+5\] move the 2y over to the left side by adding (-2y) to both sides 3y+3=-3x+5 Move 3 to the other to the right side by adding (-3) 3y=-3x+2 Divide 3 to each term y=-x+(2/3)